OpenStudy (anonymous):

write an explicit definition for the sequence -4,1,6,11...

OpenStudy (anonymous):

@mathstudent55

zepdrix (zepdrix):

Hey Belle :) Hmm this appears to be an arithmetic sequence. Each term is increasing by the same amount. How much are we increasing by each time?

OpenStudy (anonymous):

5

zepdrix (zepdrix):

Mmm k good.

zepdrix (zepdrix):

If we call the first term \(\Large\rm a_1\), then we know that \(\Large\rm a_1=-4\). And then \(\Large\rm a_2=-4+5\), yah? :o

OpenStudy (anonymous):

yeah because it equals 1 right?

zepdrix (zepdrix):

Mhm :) To get to further terms in the sequence, we add more 5's.\[\Large\rm a_3=-4+5+5\]\[\Large\rm a_3=-4+5(2)\]

zepdrix (zepdrix):

We would like to write this more `generally` though for our sequence. We want to know a `formula` for getting the "nTH" term.

OpenStudy (anonymous):

So it would be an= -4+5(n) ? the n for what number term it is?

zepdrix (zepdrix):

Notice that for the `third` term, we needed two 5's. For the `fourth` term, we would add three 5's. That pattern would repeat. So for our "nTH" term, we would start with -4 and add (n-1) 5's.

OpenStudy (anonymous):

Ooooooh, okay! I got it! Thank you so much!!

zepdrix (zepdrix):

Yay team \c:/